OpenStudy (anonymous):

how would you find the exact extreme values of this function? f(x,y) = (x-5)^(2) + (y-1)^(2) + 36 x is less than or equal to 19 y is less than or equal to 13

zepdrix (zepdrix):

Mmmmmm I don't remember.. I think for multivalued functions it's when,\[\Large\bf\sf f_x=0,\qquad\text{and}\qquad f_y=0\]right? +_+

OpenStudy (anonymous):

just differentiate it, it is implicit curve.. then you would find x=5 and y is complex number

OpenStudy (anonymous):

check Fx = 0 and Fy = 0, and also check the boundaries.

OpenStudy (anonymous):

so how would you know the max and min?

OpenStudy (anonymous):

Irrelevant: I've just come back after a few weeks, and LaTeX is back. YES!

OpenStudy (anonymous):

??

OpenStudy (whpalmer4):

So same deal as with our other problem: Expand \[f(x,y) = x^2-10x+y^2-2y+62\]Take partial derivatives and set equal to 0 \[2x-10 = 0\]\[x=5\]\[2y-2=0\]\[y=1\] Extreme value is \(f(5,1) = (5)^2-10(5)+(1)^2-2(1)+62 = 36 \) Contour plot attached